AURA-O closed today (4am Tucson time). General unrest in Chile — including Cocimbo. But looks like normality is returning — very calm in general. Italians arrived at Recinto. Spanish (TMA) to arrive soon - hopefully.
Stores open partially — long lines, also at gas stations.
Mountain is well stocked. Depending on curfew bus will go accordingly (5pm decision today — this time stick to it).
NSF & DOE visit last Saturday.
Happy with NSF sign on building.
Trip to summit: windy, but the dome moved.
All got their flights before problems started.
IT service contract was to start today as well.
IT manager still not hired — (improved) offer not accepted.
Did get an engineer, though.
Annual audit starts today.
Expect this to be handled by Daniel Calabrese.
Some murmuring on naming, but nothing concrete.
LCR-2045: commissioning descope
This shortens commissioning by 3 months, also less funding for DM (although the commissioning funding is not factored into our budget yet).
Need to look at it.
Filter carousel arrived at SLAC; filter changer and other parts may be on boat (will be this week) from France.